OpenStudy (anonymous):

Mike is three times as old as his little sister, Mary. In 4 years, he will be twice as old as Mary. How old is Mike?

OpenStudy (anonymous):

Mary = x Mike = 3x In 4 years Mike will be twice as old as mary =2(x+4)

OpenStudy (anonymous):

can u continue from there?

OpenStudy (anonymous):

No... I don't get it

OpenStudy (anonymous):

Did u follow the way i set up the variables?

OpenStudy (anonymous):

Solve a system with two equations one being x=3y and the second being y+4=2x. X-mike Y- mary

OpenStudy (anonymous):

i guess u can solve it like that too using subsitution

OpenStudy (anonymous):

I solved it differently but how were u taught to solve it in skewl?

OpenStudy (anonymous):

I soved it like this 3x+4 = 2(x+4) So basically i set up the equation descrobong mikes age in 4 yrs in two different ways and then i solved for x

OpenStudy (anonymous):

Mike is three times as old as his little sister, Mary. In 4 years, he will be twice as old as Mary. How old is Mike? Let Mike's age be "3x" yrs. Mary's age is "x" yrs. Mike's age will be "3x+4" yrs. Mary's age will be "x+4" yrs. Mike = 2(Mary) 3x+4 = 2(x+4) 3x+4 = 2x + 16 x = 4 yrs. (Mary's age NOW) 3x = 12 yrs. (Mike's age NOW) Mike is 12 years old.